I need help!!
A wholesaler is offering two different package deals of roses and carnations to florists. One package contains 20 dozen roses and 34 dozen carnations for $504.00. The other package contains 15 dozen roses and 17 dozen carnations for $327.00.
Let x = cost of one dozen roses
Let y = cost of one dozen carnations
a. Write a system of equations for this problem.
b. Which equation can be multiplied by a coefficient to create an opposite coefficient? What is that value?
c. Solve this system for the cost of one dozen roses and the cost of one dozen carnations.

Answer by stanbon(75887)

A wholesaler is offering two different package deals of roses and carnations to florists. One package contains 20 dozen roses and 34 dozen carnations for $504.00. The other package contains 15 dozen roses and 17 dozen carnations for $327.00.
Let x = cost of one dozen roses
Let y = cost of one dozen carnations
a. Write a system of equations for this problem.
1st; 20x+34y= 504
2nd: 15x+17y=327

b. Which equation can be multiplied by a coefficient to create an opposite coefficient? 2nd is the best choice
What is that value? -2
c. Solve this system for the cost of one dozen roses and the cost of one dozen carnations.
1st: 20x+34y=504
3rd: -30x-34y=-654
--------
Add 1st and 3rd to get:
-10x = -150
x = 15 (# of dozens of roses)
-------------
Substitute into 20x+34y=504 ot solve for y:
20*15+34y = 504
300 + 34y = 504
34y = 204
y = 6 (# of dozens of carnations)
===============
Cheers,
Stan H.